# Thin films spreading under the action of surface acoustic waves

### v1

@@ 4.3 Traveling wave solution
$x_{r}(t)=Ut,\qquad x_{f}(t)={\rm w}+Ut,$ (28)
where (constant) $\rm w$ is the width of the moving drop. Here, we will consider that the rear (front) contact lines of the traveling drop are given by $x_{1}^{\ast}$ ($x_{2}^{\ast}$) (see Fig. 5), since the SAW force that drives it corresponds to $h>h^{\ast}$. Then, we define $\xi=x-x_{1}^{\ast}(t)=x-Ut$ and assume that $h(x,t)=h(\xi)$, so that Eq. (21) becomes
$U\frac{dh}{d\xi}+\frac{d}{d\xi}\left[h^{3}\frac{d}{d\xi}\left(-\frac{d^{2}h}{d{\xi}^{2}}+{\rm Bo}\,h+S\phi_{s}(\xi,h)\right)\right]=0$ (29)
@@
where $J$ represents the flux. The traveling wave solution must be calculated for $0\leq\xi\leq\xi_{f}$, along with the following boundary conditions at $\xi=0$
@@
where the prime denotes $d/d\xi$. The values of $h^{\prime\prime}(0)$, $U$ and $\xi_{f}$ are determined by
@@
Note that $J$ corresponds to the flux within a thin film of thickness $h^{\ast}$ that enters the drop at $\xi_{f}$ and comes out of it at $\xi=0$ (with respect to the reference frame fixed at the drop). Also, note that $\xi_{f}$ is exactly the width of the traveling drop, $\rm w$.
We develop an iterative scheme to perform the numerical integration of Eq. (31), under the conditions given by Eq. (33). We start by guessing values of $h^{\prime\prime}(0)$, $U$ and $\xi_{f}$ to perform the integration of Eq. (31) and modify them accordingly until these conditions are satisfied within a small relative error (typically, $10^{-7}$). After obtaining convergence for a given $A$, we use these converged values of $h^{\prime\prime}(0)$, $U$ and $\xi_{f}$ as new guess values for $A+\Delta A$ (continuation). $\Delta A$ has to be very small to get a new convergence (typically, $\Delta A=10^{-5}\,$nm). Figure 11(a) shows the results of drop profiles for the values of $A$ considered so far. Clearly, drops became taller and narrower as $A$ increases. To compare the results with the experimental ones, we note that for $A=1.3$ nm, the dimensional drop speed is ${\rm v}_{f}=(\ell/t_{c})U=2.75$ mm/s, which is of the order of the experimental value. As expected, the value of $J$ is small compared to unity, $J=3.26\times 10^{-4}$. Figure 11(b) shows the dependence of drop speed, ${\rm v}_{f}$, its width, $\rm w$, and the maximum thickness, $h_{\rm max}$, on the SAW amplitude, $A$. Note that the model predicts that the speed depends approximately quadratically on the amplitude, $A$, as expected since the drop thickness increases linearly with $A$. We note that the trends shown in this figure for the travelling wave are similar to those in Fig. 10(b) obtained using full numerical solution.
@@
Figure 11: (a) Thickness profile of the traveling wave solution for: $A=0.50$ nm (green), $0.65$ nm (black), $1.04$ nm (magenta), $1.30$ nm (red), $1.69$ nm (blue). (b) Front speed, ${\rm v}_{f}$ (mm/s, black line), drop width, $\rm w$ (mm, blue line), and maximum height, $h_{\rm max}$ (mm, red line), of the traveling drop as a function of the SAW amplitude, $A$.
We note that we can use these results also to obtain an approximate value of $U$ by applying a simple formula to estimate the scale of the front speed, ${\rm v}_{f}$. By setting $J\approx 0$ and neglecting both capillary and gravity effects in Eq. (29), we have
@@
By setting $h=h_{d}/\ell$ and $\xi=2r_{d}/\ell$, this expression yields $v_{f}=U\ell/t_{c}=0.4$ mm/s for $A=1.3$ nm, which is the right order of magnitude.
5 Conclusions
This paper presents experimental results and a self-consistent model describing the spreading of a silicone oil drop under the influence of acoustic driving produced by a surface acoustic wave (SAW). Both simulations and the theoretical model show consistent features, and in particular constant spreading speed for long times. This feature of the results is specific to this type of forcing, and is different from spreading under a body force (such as gravity) where fluid thickness slowly decreases as a drop spreads, leading to decreased spreading speed (which typically scales as $h^{2}$, where $h$ is the drop thickness). Our understanding is that this feature of the results is a consequence of the particular forcing imposed by SAW, which decreases as a function of film height; therefore, in contrast to the gravity spreading, here we have a situation where the drop is essentially pushed from behind, leaving just a very thin (on the order of the precursor scale) film behind the spreading drop. The drop profile itself includes a ‘foot’ behind the main drop; this feature of the results still needs further investigation.
@@ 5 Conclusions
While not all details of the computational and experimental results are fully understood, in general, we find very reasonable agreement between the experiments and theoretical results, in particular regarding drop shapes, spreading speeds, and the dependence of the results on the SAW amplitude. For long times, both experiments and simulations show that a drop essentially translates with constant speed and without change of shape. This finding motivated us to formulate a traveling wave description of the system. The resulting solution predicts a linear increase in the drop thickness and a quadratic increase of the spreading speed with the acoustic amplitude $A$.
While many features of the experimental results have been rationalized using the formulated model, we note that there is still much to be done. In particular, further investigation is required to understand all the features of the calculated drop profiles, and also three-dimensional aspects of the results remain to be simulated and understood. Consideration of such effects will be the subject of future work.
@@ Appendix A Calibration of applied voltage versus SAW displacement
We use the normal displacement amplitude (particle displacement of the SAW), $A$, to represent the SAW strength. In the experiments, we control this amplitude by changing the applied voltage of the signal generator. Therefore, we conducted an experiment to obtain a calibration curve between the applied voltage and amplitude, as shown in Fig 12. We measure $A$ over a surface of $1\times 1~{}\text{mm}^{2}$ about 2 mm away from the SAW actuator using laser doppler vibrometer. We find an approximately linear dependence of $A$ on voltage, consistently with the literature (Ballantine Jr et al., 1996).
@@ Appendix B Computational implementation
At the domain ends, we apply Dirichlet boundary conditions $\tilde{h}=\tilde{h}_{p}$, where $\tilde{h}_{p}$ is the precursor film thickness, and $\partial h/\partial x=0$. In order to achieve convergent numerical simulations, we discretize the domain with steps $\Delta x=h_{p}$.

### v2

@@ IV.3 Traveling wave solution
$x_{\rm r}(t)=Ut,\qquad x_{\rm f}(t)={\rm w}+Ut,$ (28)
where (constant) $\rm w$ is the width of the moving drop. Here, we will consider that the rear (front) contact lines of the traveling drop are given by $x_{1}^{\ast}$ ($x_{2}^{\ast}$) (see Fig. 5), since the SAW force that drives it is felt only where $h>h^{\ast}$. Then, we define $\xi=x-x_{1}^{\ast}(t)=x-Ut$ and assume that $h(x,t)=h(\xi)$, so that Eq. (21) becomes (see also [57, 58, 59] for similar studies without SAW forces)
$U\frac{dh}{d\xi}+\frac{d}{d\xi}\left[h^{3}\frac{d}{d\xi}\left(-\frac{d^{2}h}{d{\xi}^{2}}+{\rm Bo}\,h+S\phi_{s}(\xi,h)\right)\right]=0,$ (29)
@@
where $J$ represents the flux. The traveling wave solution must be calculated for $0\leq\xi\leq\xi_{\rm f}$, along with the following boundary conditions at $\xi=0$
@@
where the prime denotes $d/d\xi$. The values of $h^{\prime\prime}(0)$, $U$ and $\xi_{\rm f}$ are determined by
@@
Note that $J$ corresponds to the flux within a thin film of thickness $h^{\ast}$ that enters the drop at $\xi_{\rm f}$ and comes out of it at $\xi=0$ (with respect to the reference frame fixed at the drop). Also, note that $\xi_{\rm f}$ is exactly the width of the traveling drop, $\rm w$.
We developed an iterative scheme to perform the numerical integration of Eq. (31), under the conditions given by Eq. (33). We start by guessing values of $h^{\prime\prime}(0)$, $U$ and $\xi_{\rm f}$ to perform the integration of Eq. (31) and modify them accordingly until these conditions are satisfied within a small relative error (typically, $10^{-7}$). After obtaining convergence for a given $A$, we use these converged values of $h^{\prime\prime}(0)$, $U$ and $\xi_{\rm f}$ as new guess values for $A+\Delta A$ (continuation). $\Delta A$ has to be very small to reach convergence (typically, $\Delta A=10^{-5}\,$nm). Figure 11(a) shows the drop profiles for the values of $A$ considered so far. Clearly, drops became taller and narrower as $A$ increases. To compare the results with the experimental ones, we note that for $A=1.3$ nm, the dimensional drop speed is ${\rm v}_{\rm f}=(\ell/t_{c})U=2.75$ mm/s, which is of the order of the experimental value. As expected, the value of $J$ is small compared to unity, $J=3.26\times 10^{-4}$. Figure 11(b) shows the dependence of drop speed, ${\rm v}_{\rm f}$, its width, $\rm w$, and the maximum thickness, $h_{\rm max}$, on $A$. Note that the model predicts that the speed depends approximately quadratically on the amplitude, $A$, as expected since the drop thickness increases linearly with $A$. We note that the trends shown in this figure for the traveling wave are similar to those in Fig. 10(b) obtained using the full numerical solution.
@@
Figure 11: (a) Thickness profile of the traveling wave solution for: $A=0.50$ nm (green), $0.65$ nm (black), $1.04$ nm (magenta), $1.30$ nm (red), $1.69$ nm (blue). (b) Front speed, ${\rm v}_{\rm f}$ (mm/s, black line), drop width, $\rm w$ (mm, blue line), and maximum height, $h_{\rm max}$ (mm, red line), of the traveling drop as a function of $A$.
We note that we can use these results also to obtain an approximate value of $U$ by applying a simple formula to estimate the front speed, ${\rm v}_{\rm f}$. By setting $J\approx 0$ and neglecting both capillary and gravity effects in Eq. (29), we have
@@
By setting $h=h_{d}/\ell$ and $\xi=2r_{d}/\ell$, this expression yields ${\rm v_{f}}=U\ell/t_{c}=0.4$ mm/s for $A=1.3$ nm, which is the right order of magnitude.
V Conclusions
In this work, we consider experimentally, theoretically, and computationally the problem of a spreading silicone oil drop driven by a surface acoustic wave (SAW) propagating in the supporting substrate. Our particular focus is on drops of millimetric thickness, for which the main driving mechanism is the Eckart streaming. Therefore, we are considering a different regime from that relevant for much thinner drops, for which Schlichting and Rayleigh streaming effects are dominant. To the best of our knowledge, modeling the dynamics of millimetric drops under the action of SAW and capillary stresses has not been considered so far in the literature.
While developing our theoretical description of drop spreading, it became apparent that a crucial effect, which requires careful modeling, involves the attenuation of the SAW under the evolving drop. The model we derived, assuming the thin film limit and implementing the long-wave approximation, shows that the drop is essentially pushed by the SAW forcing from behind, in contrast to situations where flow is driven by a body force such as gravity (for example, spreading on an inclined plane). A consequence of the driving mechanism is that the dynamics itself is substantially different from the familiar case of drops driven by gravity, where the volume force is a constant and not rapidly decreasing along the drop, as it occurs with the SAW force. In particular, we find that the drops tend to keep almost the same shape with time and spreading progress, and also that they reach essentially constant spreading speed at long times. This contrasts with the gravity-driven case, where the spreading speed decreases due to drop thinning. The main features of our theoretical and computational results are in reasonable qualitative and sometimes even quantitative agreement with the experiments, supporting further the basic premises leading to the developed acousto-fluidic model.
@@ V Conclusions
While not all details of the experimental results are fully understood, in general, we find very reasonable agreement between the experiments and theoretical results, in particular regarding drop shapes, spreading speeds, and the dependence of the results on the SAW displacement amplitude at the solid surface (intensity). Both theoretical and experimental results show asymmetric drop shapes with a thin film left behind, and both appear to be reaching a constant spreading speed asymptotically in time. The latter finding motivated us to formulate a simplified traveling wave model. The resulting analytical solution predicts a linear increase in the drop thickness and a quadratic increase of the spreading speed with the acoustic intensity (SAW normal displacement amplitude at the solid surface), $A$. Such predicted trends are consistent with the predictions of the full theoretical model and the experiments.
While many features of the experimental results have been rationalized using our theoretical models and simulations, much remains to be done. Further investigation is required to shed light on the features of dynamic drop profiles under SAW excitation. The transition from three-dimensional (3D) to quasi-two-dimensional (2D) dynamics is also intriguing: under strong SAW excitation, a drop spreads along the path of the SAW and may be described to leading order using the 2D model given here. However, at weak SAW excitation, where the drop also spreads radially due to the relaxation of capillary stress at the free drop surface, the drop dynamics is inherently of the 3D type. Hence, 3D aspects of drop dynamics, in particular in cases where the contributions to flow from capillary and SAW stresses are comparable, remain to be simulated and understood. Furthermore, the main physical mechanisms involved in the formation of the trailing film behind the spreading drops should be understood in more detail.
@@ Appendix A Calibration of applied voltage versus SAW displacement
We use the normal displacement amplitude (particle displacement of the SAW) at the solid surface, $A$, to represent the SAW strength (intensity). In the experiments, we control this amplitude by changing the applied voltage of the signal generator. Therefore, we conducted an experiment to obtain a calibration curve between the applied voltage and amplitude, as shown in Fig 12. We measure $A$ over a surface of $1\times 1~{}\text{mm}^{2}$ about 2 mm away from the SAW actuator using laser doppler vibrometer. We find an approximately linear dependence of $A$ on voltage, consistently with the literature [60].
@@ Appendix B Computational implementation
At the domain ends, we apply Dirichlet boundary conditions $\tilde{h}=\tilde{h}_{p}$, where $\tilde{h}_{p}$ is the precursor film thickness, and $\partial h/\partial x=0$. In order to achieve convergent numerical simulations, we discretize the domain using $\Delta x=h_{p}$.